# Quantum accessible information and classical entropy inequalities

### v2

@@ 4.5 The case of flat pyramid
Arguing as in the previous section, we suggest that the hypothetical optimal observable is given by (61) with $t=1$ for $m\geq 7$ and $t=\infty$ for $m\leq 6$. The corresponding tight entropy inequality is
$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\left\{\begin{array}[]{cc}1,&m\leq 6;\\ \log m-\frac{m-2}{m}\log(m-1),&m\geq 7,\end{array}\right.$ (75)
for complex $z_{j}$ satisfying (74). Indeed, for $m\leq 6$ we can use the argument leading to (72) with $\sum_{j=1}^{m}z_{j}=0.$ On the other hand, for $m\geq 7$ we should use (66) with $\sum_{j=1}^{m}z_{j}=0$ and $p=\frac{m-1}{m}.$ Then the first term in the righthand side vanish, while $\tilde{\mu}_{1}(p)$ gets the value
$-\frac{m-1}{m}\log\frac{m-1}{m}-\frac{1}{m}\log\frac{1}{m(m-1)}=\log m-\frac{m-2}{m}\log(m-1),$ (76)
which is greater than 1 for $m\leq 6$, but less for $m\geq 7$.
@@
Proof.
By lemma 1 below we can restrict to the case of real $z_{j}$. Introducing the Lagrange function
$L\left(z_{1},\dots,z_{m}\right)=-\sum_{j=1}^{m}z_{j}^{2}\log z_{j}^{2}-\lambda\sum_{j=1}^{m}z_{j}^{2}-\mu\sum_{j=1}^{m}z_{j},$
which is continuously differentiable, we have the necessary condition for a minimizer of the lefthand side of (75)
$-\frac{\partial L}{\partial z_{j}}=2z_{j}\log ez_{j}^{2}+2\lambda z_{j}+\mu=0,\quad j=1,\dots,m.$
Case 1. The minimizer has a zero component $z_{j}$. Then $\mu=0$ and the equation $02z\log ez^{2}+2\lambda z=$ 0 has the roots $0,c,-c.$ Therefore the minimizer should have the form $(\underset{k\text{ times}}{\underbrace{c,-c,\dots,c,-c,}}\,0,\dots,0),$ where $12kc^{2}=$ 1 giving the value $-2kc^{2}\log c^{2}=\log 2k$ which has the minimum equal to 1 for $k=1.$
Case 2. The minimizer has no zero components. The equation $02z\log ez^{2}+2\lambda z+\mu=$ 0 has at most three real, nonzero roots $c_{1},c_{2},c_{3}$, which enter the minimizer with the multiplicities $k_{1},k_{2},$ $k_{3},$ such that $k_{1}+k_{2}+k_{3}=m.$ Then consider the Lagrange function
$\tilde{L}\left(c_{1},c_{2},c_{3};k_{1},k_{2},k_{3}\right)=-\sum_{j=1}^{3}k_{j}c_{j}^{2}\log c_{j}^{2}-\lambda\sum_{j=1}^{3}k_{j}c_{j}^{2}-\mu\sum_{j=1}^{3}k_{j}c_{j}-\nu\sum_{j=1}^{3}k_{j}.$
The necessary condition is
$-\frac{\partial\tilde{L}}{\partial c_{j}}=2k_{j}c_{j}\log ec_{j}^{2}+2\lambda k_{j}c_{j}+\mu k_{j}=0,$
$-\frac{\partial\tilde{L}}{\partial k_{j}}=c_{j}^{2}\log ec_{j}^{2}+\lambda c_{j}^{2}+\mu c_{j}+\nu=0,\quad j=1,2,3.$
Taking into account that $c_{j}\neq 0,k_{j}\geq 1,$ we can multiply the first equation by $c_{j}/\left(2k_{j}\right)$ without lost of information and substract the second equation, resulting in
$c_{j}^{2}-\left(\mu/2\right)c_{j}-\nu=0,$
which has at most two different roots, say $c_{1},c_{2}$, entering the minimizer with the multiplicities $k_{1},k_{2},$ such that $k_{1}+k_{2}=m.$ Then $k_{1}c_{1}+k_{2}c_{2}=0,k_{1}c_{1}^{2}+k_{2}c_{2}^{2}=1$ whence $c_{1}=\sqrt{\frac{k_{2}}{k_{1}m}},c_{1}=-\sqrt{\frac{k_{1}}{k_{2}m}}$ and the value of the lefthand side in (75) is
$-k_{1}c_{1}^{2}\log c_{1}^{2}-k_{2}c_{2}^{2}\log c_{2}^{2}=-\frac{k_{2}}{m}\log\frac{k_{2}}{k_{1}m}-\frac{k_{1}}{m}\log\frac{k_{1}}{k_{2}m}.$
This is a concave function of $k_{1}=m-k_{2}$ symmetric on the interval $[1,m-1]$ hence it attains the minimum at the edges of the interval, which is equal to (76). Since (76) is greater than 1 for $m\leq 6$, but less for $m\geq 7$, we arrive to (75). ∎
@@
Lemma 1.
If the inequality (75) holds for all real $z_{j}$ satisfying the conditions (74), then it holds for all complex $z_{j}$ satisfying the same conditions.
Proof.
Let $z_{j}=x_{j}+iy_{j}$ for real $x_{j},y_{j}$, then (74) amounts to
$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1,\quad\sum_{j=1}^{m}x_{j}=0,\,\sum_{j=1}^{m}y_{j}=0.$ (77)
[…]
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}$ $=$ $\sum_{j=1}^{m}f\left(P\frac{x_{j}^{2}}{P}+(1-P)\frac{y_{j}^{2}}{1-P}\right)$
$\leq$ $P\sum_{j=1}^{m}f\left(\frac{x_{j}^{2}}{P}\right)+(1-P)\sum_{j=1}^{m}f\left(\frac{y_{j}^{2}}{1-P}\right)$
$=$ $P\sum_{j=1}^{m}\frac{x_{j}^{2}}{P}\log\left(\frac{x_{j}^{2}}{P}\right)$
[…]
because
$\sum_{j=1}^{m}\frac{x_{j}^{2}}{P}=\sum_{j=1}^{m}\frac{y_{j}^{2}}{1-P}=1,\quad\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}=\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}=0,$
and by assumption (75) holds for real $z_{j}$. ∎
5 Proofs
5.1 Proof of theorem 2
@@
$F(t):=H(t)-\Bigl{(}\mu_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl{)}^{2}-\mu_{1}(p)\Bigl{)}\geq 0,$ (78)
@@
In the special case $p=1$, the inequality (78) is trivially satisfied by substituting $\mu_{0}(1):=0,\ \mu_{1}(1)=0$.
@@
for the boundary point $t_{*}$ and the transversal direction $a$ has the derivative $\infty$- ∞ at zero:
@@
up to permutations, for some $1\leq k\leq m$ and $q\in[0,1]$ (for $k=m$ it is assumed that $q=1/m$). One of the critical points is equal to $(p,(1-p)/(m-1),...,(1-p)/(m-1))$, in which the value of the function is equal to 00.
@@
Lemma 3.
Let $A>0,\ B,\ C\in\mathbb{R}$. Then the function
$f(t)=h(t)-(A\sqrt{t(1-t)}+Bt+C)$ (82)
[…]
Proof.
$f^{\prime}(0)=-\infty,\quad f^{\prime}(1)=+\infty,$
therefore local minima are on the interval $(0,1)$.
In the calculations we will move on to the natural logarithm.
$f^{\prime}(t)=-\ln\dfrac{t}{1-t}-\dfrac{A\sqrt{t(1-t)}}{2}\Bigl{(}\dfrac{1}{t}-\dfrac{1}{1-t}\Bigl{)}-B=$
$=-\ln\dfrac{t}{1-t}-\dfrac{A(1-2t)}{2\sqrt{t(1-t)}}-B;$
$f^{\prime\prime}(t)=-\dfrac{1}{t(1-t)}+\dfrac{A}{4(t(1-t))^{3/2}}.$
$f^{\prime\prime}(t)=0\quad\Leftrightarrow\quad A=4\sqrt{t(1-t)}.$
[…]
@@
$\alpha\in{(0,1/2],\quad p\in[1-\alpha,1},\quad\beta\in[\alpha,1/2],\quad t\in[0,1].$
@@
$\mu_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\mu_{1}(p,\alpha)$ (83)
@@
for which $A=2\mu_{0}(p,\alpha)\sqrt{\beta(1-\beta)}$, $B=\log\dfrac{1-\beta}{\beta}+\mu_{0}(p,\alpha)(2\beta-1)$ in the form (82).
@@
    In this case there is minimum point $t_{*}=\beta$ ($f^{\prime}(\beta)=0$) of the function $f(t)$ on $[0,1/2]$ due to the Lemma 3. One has
@@
    Let $L(p)=\sqrt{\alpha p}\ln\dfrac{p}{\alpha}+\sqrt{(1-\alpha)(1-p)}\ln\dfrac{1-p}{1-\alpha}$, then
@@
      $(d/dp)^{2}L(p)=-\dfrac{\sqrt{\alpha}\ln(p/\alpha)}{4p^{3/2}}-\dfrac{\sqrt{1-\alpha}\ln((1-p)/(1-\alpha))}{4(1-p)^{3/2}}\geq 0$
@@
      $L(1-\alpha)=0$ and
      $L^{\prime}(1-\alpha)=\sqrt{\dfrac{\alpha}{1-\alpha}}-1+\dfrac{1}{2}\ln\dfrac{\alpha}{1-\alpha}+\dfrac{1}{2}\sqrt{\dfrac{\alpha}{1-\alpha}}\ln\dfrac{1-\alpha}{\alpha}\geq 0\quad\Leftrightarrow$
@@
    Therefore, $L(p)\geq 0$ and the inequality (85) is satisfied.
@@
    This condition implies that the function $f(t)$ has the minimum point $t_{*}$ on $[1/2,1]$. This point is characterized by the equation
@@
    The value of $t_{*}$ depends on $\beta$, and
@@
    ×$\times\Bigl{(}\sqrt{\beta t_{*}}+\sqrt{(1-\beta)(1-t_{*})}\Bigl{)}.$
@@
    $\sqrt{\dfrac{1-\beta}{\beta}\dfrac{t_{*}}{1-t_{*}}}-\sqrt{\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}}+\ln\Bigl{(}\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}\Bigl{)}\equiv C-1/C-2\ln C\geq 0$
    for $C\geq 1$.
@@
5.2 Proof of theorem 3
This section is devoted to proving the inequality (66) which verificates the condition (ii.a) in case of obtuse pyramid. Namely, if non-negative parameter $p<1$ is larger than the maximum of $\dfrac{m-1}{m}$ and $p(m)$ that is defined in (71), then for complex numbers $z_{1},...,z_{m}$ forming probability distribution $(|z_{1}|^{2},...,|z_{m}|^{2})$, the inequality
$-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}\geq\tilde{\mu}_{0}(p)\Bigl{|}\sum\limits_{j=1}^{m}z_{j}\Bigl{|}^{2}-\tilde{\mu}_{1}(p).$
@@ 5.2 Proof of theorem 3
For $p=1$, which corresponds to the orthonormal case, the inequality reduces to a trivial one: $H(|z_{1}|^{2},...,|z_{m}|^{2})\geq 0$.
Proof.
Denote $z_{j}=x_{j}+iy_{j}$ and consider the Lagrange function
$L(x_{1},...,x_{m},y_{1},...,y_{m})=-\sum\limits_{k=1}^{m}(x_{k}^{2}+y_{k}^{2})\log(x_{k}^{2}+y_{k}^{2})-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}x_{k}\Bigl{)}^{2}-$ (86)
$-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}y_{k}\Bigl{)}^{2}+\tilde{\mu}_{1}(p)-\lambda\Bigl{(}\sum\limits_{k=1}^{m}(x_{k}^{2}+y_{k}^{2})-1\Bigl{)}.$ (87)
The necessary conditions are:
$\begin{cases}-\dfrac{1}{2}\dfrac{\partial L}{\partial x_{k}}=x_{k}\Bigl{(}\log(x_{k}^{2}+y_{k}^{2})+\log e\Bigl{)}+\tilde{\mu}_{0}(p)\sum\limits_{k=1}^{m}x_{j}+\lambda x_{k},\\ -\dfrac{1}{2}\dfrac{\partial L}{\partial y_{k}}=y_{k}\Bigl{(}\log(x_{k}^{2}+y_{k}^{2})+\log e\Bigl{)}+\tilde{\mu}_{0}(p)\sum\limits_{k=1}^{m}y_{j}+\lambda y_{k}.\end{cases}$ (88)
for all $1\leq k\leq m$.
Hence, we get
$\dfrac{1}{2}\Bigl{(}y_{k}\dfrac{\partial L}{\partial x_{k}}-x_{k}\dfrac{\partial L}{\partial y_{k}}\Bigl{)}=\tilde{\mu}_{0}(p)\Bigl{(}x_{k}\sum\limits_{j=1}^{m}y_{j}-y_{k}\sum\limits_{j=1}^{m}x_{j}\Bigl{)}=0.$ (89)
Case 1. $\sum\limits_{j=1}^{m}z_{j}=0$ or $\tilde{\mu}_{0}(p)=0$.
We get the inequality $-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}\geq-\tilde{\mu}_{1}(p)$. Taking into account proposition 2 for the flat pyramid, it is sufficient to prove the lemma
Lemma 5.
Let $m\geq 2$ be a natural number.
  1. 1.
    If $\dfrac{m-1}{m}\leq p<1$, then $-\tilde{\mu}_{1}(p)\leq-\dfrac{m-2}{m}\log(m-1)+\log m$;
  2. 2.
    if $m\leq 6$ and $p(m)\leq p<1$, then $-\tilde{\mu}_{1}(p)\leq 1$.
Case 2. $\sum\limits_{j=1}^{m}z_{j}\neq 0$. Then the condition (89) reduces to proportionality of two vectors $(x_{1},...,x_{m})$ and $(y_{1},...,y_{m})$. Due to the symmetry of (86) under multiplication by a scalar $e^{i\varphi},\ \varphi\in\mathbb{R}$, the problem reduces to the case of real variables $\{z_{j}\}$.
@@
For the real variables $\{z_{j}\}$ let us consider the new Lagrange function
$L_{(1)}(z_{1},...,z_{k})=-\sum\limits_{k=1}^{m}(z_{k}^{2})\log(z_{k}^{2})-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}z_{k}\Bigl{)}^{2}+\tilde{\mu}_{1}(p)-\lambda\Bigl{(}\sum\limits_{k=1}^{m}z_{k}^{2}-1\Bigl{)}.$
Thus, the necessary condition is
$-\dfrac{1}{2}\dfrac{\partial L_{(1)}}{\partial z_{k}}=z_{k}(\log z_{k}^{2}+\log e)+\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}z_{j}\Bigl{)}+\lambda z_{k}=0,\quad 1\leq k\leq m,$
@@
and the set $\{z_{j}\}$ has at most $3$ different values $c_{1},c_{2},c_{3}$ with multiplicities $k_{1},k_{k},k_{3}$, since the equation
$z(\log z^{2}+\log e)+\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}z_{j}\Bigl{)}+\lambda z=0$
@@
for a variable $z$ has at most $3$ real solutions. The fact the the equation of the form $z\log z^{2}+az+b=0$ for real $a$ and $b$ has no more then $3$ real roots was actually used in proof of proposition 2. To reduce the number of different values to two, we will introduce a new Lagrange function $L_{(2)}$. The critical points of $L_{(1)}$ are related with critical points of the Lagrange function
$L_{(2)}(c_{1},c_{2},c_{3},k_{1},k_{2},k_{3})=-\sum\limits_{j=1}^{3}k_{j}c_{j}^{2}\log c_{j}^{2}-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}k_{j}c_{j}\Bigl{)}^{2}-$
$-\lambda\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}^{2}-1\Bigl{)}-\nu\Bigl{(}\sum\limits_{j=1}^{3}k_{j}-m\Bigl{)}.$
@@
$-\dfrac{1}{2}\dfrac{\partial L_{(2)}}{\partial c_{l}}=k_{l}c_{l}(\log c_{l}^{2}+\log e)+\tilde{\mu}_{0}(p)k_{l}\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}+\lambda k_{l}c_{l}=0,$
$-\dfrac{\partial L_{(2)}}{\partial k_{l}}=c_{l}^{2}\log c_{l}^{2}+2\tilde{\mu}_{0}(p)c_{l}\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}+\lambda c_{l}^{2}+\nu=0$
@@
By transformation, we have
$-\dfrac{1}{2}c_{l}\dfrac{\partial L_{(2)}}{\partial c_{l}}+k_{l}\dfrac{\partial L_{(2)}}{\partial k_{l}}=k_{l}c_{l}\Bigl{(}c_{l}\log e-\tilde{\mu}_{0}(p)\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}-\nu k_{l}=0.$
Consider the next subcases.
@@
For all $1\leq l\leq 3$ it holds $c_{l}^{2}\log e=\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}+\nu$, which can take place, if the set $\{c_{1},c_{2},c_{3}\}$ has no more than $2$ different values. This can be regarded as the next case.
@@
Up to permutation, $(z_{j})_{j=1}^{m}=(\underbrace{c_{1},...,c_{1}}_{k_{1}\text{ times}},\underbrace{c_{2},...,c_{2}}_{k_{2}\text{ times}})$, $c_{1}\geq 0,c_{2}\in\mathbb{R}$.

### v3

@@ 4.5 The case of flat pyramid
Arguing as in the previous section, we suggest that the optimal observable is given by (61) with $t=1$ for $m\geq 7$ and $t=\infty$ for $m\leq 6$. The corresponding tight entropy inequality is
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\left\{\begin{array}[]{cc}1,&m\leq 6;\\ \log m-\frac{m-2}{m}\log(m-1),&m\geq 7,\end{array}\right$. (82)
for complex $z_{j}$ satisfying (81)  (due to the second statement of lemma 1 one can restrict here to real $z_{j}$ ). Indeed, for $m\leq 6$ we can use the argument leading to inequality (79) with $\sum_{j=1}^{m}z_{j}=0$ and $p=p(m)$. Then (78) implies the first line in the right-hand side of (82). On the other hand, for $m\geq 7$ we should use the inequality (67) with $\sum_{j=1}^{m}z_{j}=0$ and $p=\frac{m-1}{m}.$ Then the first term in the righthand side vanish, while $\tilde{\mu}_{1}(p)\$gets the value (74) i.e. the second line in the righthand side of (82), which is greater than 1 for $m\leq 6$, but less for $m\geq 7$. We thus arrive at the following special case of theorem 3:
@@
5 Suggestions for proofs
5.1 The case of theorem 2
@@
$F(t)\equiv H(t)-\Bigl{(}\mu_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl{)}^{2}-\mu_{1}(p)\Bigl{)}\geq 0,$ (83)
@@
In the special case $p=1$, the inequality (83) is trivially satisfied by substituting $\mu_{0}(1)\equiv 0,\ \mu_{1}(1)=0$.
@@
for the boundary point $t_{*}$ and the transversal direction $a$ has the derivative -$\infty$ at zero:
@@
up to permutations, for some $1\leq k\leq m$ and $q\in[0,1]$ (for $k=m$ it is assumed that $q=1/m$). One of the critical points is equal to (p,(1p)/(m1),,(1p)/(m1))(p,(1-p)/(m-1),...,(1-p)/(m-1)), in which the value of the function is equal to 0.
@@
The proof of the following lemma is based on the behavior of the difference between binary entropy and a function containing both a linear component and a square root term $\sqrt{t(1-t)}$. We conduct a qualitative analysis of the minimizer of this function. The global minimum (there can be no more than two local minima) appears on either the interval $[0,1/2]$ or $[1/2,1]$. Importantly, this segment contains no other critical points. Therefore, minimizing the function reduces to identifying the critical point within the appropriate interval and evaluating the function at that point.
A detailed proof of this statement will be provided at the end of the section.
@@
$\alpha\in{(0,1/2],\quad p\in[1-\alpha,1)},\quad\beta\in[\alpha,1/2],\quad t\in[0,1].$
@@
$\mu_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\mu_{1}(p,\alpha)$
@@
for which $A=2\mu_{0}(p,\alpha)\sqrt{\beta(1-\beta)}$, $B=\log\dfrac{1-\beta}{\beta}+\mu_{0}(p,\alpha)(2\beta-1)$ in the form (89) (see lemma 4 below).
@@
    In this case there is minimum point $t_{*}=\beta$ ($f^{\prime}(\beta)=0$) of the function $f(t)$ on $[0,1/2]$ due to the lemma 4 presented later. One has
@@
    Let $V(p)=\sqrt{\alpha p}\ln\dfrac{p}{\alpha}+\sqrt{(1-\alpha)(1-p)}\ln\dfrac{1-p}{1-\alpha}$, then
@@
      $(d/dp)^{2}V(p)=-\dfrac{\sqrt{\alpha}\ln(p/\alpha)}{4p^{3/2}}-\dfrac{\sqrt{1-\alpha}\ln((1-p)/(1-\alpha))}{4(1-p)^{3/2}}\geq 0$
@@
      $V(1-\alpha)=0$ and
      $V^{\prime}(1-\alpha)=\sqrt{\dfrac{\alpha}{1-\alpha}}-1+\dfrac{1}{2}\ln\dfrac{\alpha}{1-\alpha}+\dfrac{1}{2}\sqrt{\dfrac{\alpha}{1-\alpha}}\ln\dfrac{1-\alpha}{\alpha}\geq 0\quad\Leftrightarrow$
@@
    Therefore, $V(p)\geq 0$ and the inequality (88) is satisfied.
@@
    This condition implies that the function $f(t)$ has the minimum point $t_{*}$ on $[1/2,1]$ (lemma 4 is applied in the same way). This point is characterized by the equation
@@
    The value of $t_{*}$ depends on $\beta$, and for the function $f(\beta,t)\equiv f(t)$ it holds $\dfrac{d}{d\beta}f(\beta,t_{*}(\beta))=\dfrac{\partial}{\partial\beta}f(\beta,t_{*}(\beta))+\dfrac{dt_{*}(\beta)}{d\beta}\dfrac{\partial}{\partial t}f(\beta,t_{*}(\beta))=\dfrac{\partial}{\partial\beta}f(\beta,t_{*}(\beta))$ since $t_{*}$ is critical for $f(t)$. Then,
@@
    $\times\Bigl{(}\sqrt{\beta t_{*}}+\sqrt{(1-\beta)(1-t_{*})}\Bigl{)}.$
@@
    $\sqrt{\dfrac{1-\beta}{\beta}\dfrac{t_{*}}{1-t_{*}}}-\sqrt{\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}}+\ln\Bigl{(}\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}\Bigl{)}=R-1/R-2\ln R\geq 0$
    where $R\equiv\sqrt{\dfrac{1-\beta}{\beta}\dfrac{t_{*}}{1-t_{*}}}\geq 1$.
@@
The proof of lemma 3 is based on a simple fact of analysis:
Lemma 4.
Let $A>0,\ B,\ C\in\mathbb{R}$ (The notation of the coefficients is used only within the framework of the formulation of the lemma and its proof). Then the function
$g(t)=h(t)-(A\sqrt{t(1-t)}+Bt+C)$ (89)
[…]
Proof.
$g^{\prime}(0)=-\infty,\quad g^{\prime}(1)=+\infty,$
therefore local minima are on the interval $(0,1)$.
In the calculations we will move on to the natural logarithm.
$g^{\prime}(t)=-\ln\dfrac{t}{1-t}-\dfrac{A\sqrt{t(1-t)}}{2}\Bigl{(}\dfrac{1}{t}-\dfrac{1}{1-t}\Bigl{)}-B=$
$=-\ln\dfrac{t}{1-t}-\dfrac{A(1-2t)}{2\sqrt{t(1-t)}}-B;$
$g^{\prime\prime}(t)=-\dfrac{1}{t(1-t)}+\dfrac{A}{4(t(1-t))^{3/2}}.$
$g^{\prime\prime}(t)=0\quad\Leftrightarrow\quad A=4\sqrt{t(1-t)}.$
The last statement obviously follows from the symmetric case. Indeed, for $B<0$, it holds $g(t)\leq g(1-t)$ for $t\in[1/2,1]$, therefore the global minimum lies on the segment $[1/2,1]$. Let s, 1ss,\ 1-s ($0<s<1/2$) be zeros of $g^{\prime\prime}(t)$. On the segment [s,1s][s,1-s] the function $g(t)^{\prime}$ is decreasing, on the segments [0,s],[1s,1][0,s],\ [1-s,1] the function $g(t)^{\prime}$ increases, and due to symmetry $(g(t)+Bt)^{\prime}$ has root $1/2$. Therefore, on the segment $[1/2,1]$ when $B<0$ the function $g(t)$ has no more than one critical point. ∎
5.2 The case of theorem 3
This section is devoted to proving the inequality (67) which verifies the condition (ii.a) in the case of obtuse pyramid. Namely, if the parameter $p<1$ satisfies $p\geq\max\left\{\dfrac{m-1}{m},p(m)\right\}$ with $p(m)$ defined in (77), then for complex numbers $z_{1},...,z_{m}$ satisfying (66) the inequality
$\tilde{F}(z)\equiv-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}-\tilde{\mu}_{0}(p)\Bigl{|}\sum\limits_{j=1}^{m}z_{j}\Bigl{|}^{2}+\tilde{\mu}_{1}(p)\geq 0.$ (90)
@@ 5.2 The case of theorem 3
Suggestion for proof.
As it has been established in lemma 1, a real case is enough for analysis.
@@
For real vectors $(z_{j})_{j=1}^{m}$ from the unit sphere let us consider the Lagrange function
$L(z_{1},...,z_{k})=-\sum\limits_{k=1}^{m}z_{k}^{2}\log z_{k}^{2}-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}z_{k}\Bigl{)}^{2}+\tilde{\mu}_{1}(p)-\lambda\Bigl{(}\sum\limits_{k=1}^{m}z_{k}^{2}-1\Bigl{)}.$
The necessary condition is
$\dfrac{1}{2}\dfrac{\partial L}{\partial z_{k}}=z_{k}\log z_{k}+\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}z_{j}\Bigl{)}+(\lambda+\log e)z_{k}=0,\quad 1\leq k\leq m,$
@@
and the set $\{z_{j}\}$ has at most three different values $s_{1},s_{2},s_{3}$ with multiplicities $k_{1},k_{2},k_{3}\geq 0$ ($k_{1}+k_{2}+k_{3}=m$), since the equation
$z\log z^{2}+\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}z_{j}\Bigl{)}+(\lambda+\log e)z=0$ (91)
@@
for a variable $z$ has at most three real solutions. The fact the equation of the form $z\log z^{2}+az+b=0$ for real $a$ and $b$ has no more then three real roots was actually used in proof of lemma 2 for equation (86) which is similar to (91) with $\leq 2$ roots. However, here we consider the entire real line, so the maximum number of real roots is three.
Consider the following cases.
Case 1. The minimizer has a zero component.
Then $\lambda+\log e=0$ and the equation (91) has roots 0 and $\pm s$. Therefore the minimizer of the function $\tilde{F}(z)$ should have the form $(\underbrace{s,-s,...,s,-s}_{k\text{ times}},0,...,0)$, where $2ks^{2}=1$ giving the value $2ks^{2}\log s^{2}+\tilde{\mu}_{1}(p)=\log 2k+\tilde{\mu}_{1}(p)$ which has the minimum equal to $1+\tilde{\mu}_{1}(p)$ for $k=1$. The value $1+\tilde{\mu}_{1}(p)$ is nonnegative for the following reason.
Thus for $\frac{m-1}{m}\leq p<1$ we have
$\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}\Bigl{(}\frac{m-1}{m}\Bigl{)}=-\dfrac{m-2}{m}\log(m-1)+\log m$
due to (74). For $p(m)\leq p<1$ we have
$\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}(p(m))=1,$
because of (78). Therefore, $1+\tilde{\mu}_{1}(p)\geq 0$ for $p>\max(p(m),(m-1)/m)$. Note that $\tilde{\mu}_{1}(p)\geq\tilde{\mu}_{1}((m-1)/m)\geq\tilde{\mu}_{1}(p(m))=-1$ if p(m)<(m1)/mp(m)<(m-1)/m.
Case 2. The minimizer has no zero components.
To reduce the number of different values to two, we will introduce a new Lagrange function $\hat{L}$. The critical points of $L$ are related with critical points of the Lagrange function
$\hat{L}(s_{1},s_{2},s_{3},k_{1},k_{2},k_{3})=-\sum\limits_{j=1}^{3}k_{j}s_{j}^{2}\log s_{j}^{2}-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}k_{j}s_{j}\Bigl{)}^{2}$-
$-\lambda\Bigl{(}\sum\limits_{j=1}^{3}k_{j}s_{j}^{2}-1\Bigl{)}-\nu\Bigl{(}\sum\limits_{j=1}^{3}k_{j}-m\Bigl{)}.$
@@
$-\dfrac{1}{2}\dfrac{\partial\hat{L}}{\partial s_{l}}=k_{l}s_{l}\log s_{l}^{2}+\tilde{\mu}_{0}(p)k_{l}\Bigl{(}\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl{)}+(\lambda+\log e)k_{l}s_{l}=0,$
$-\dfrac{\partial\hat{L}}{\partial k_{l}}=s_{l}^{2}\log s_{l}^{2}+2\tilde{\mu}_{0}(p)s_{l}\Bigl{(}\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl{)}+\lambda s_{l}^{2}+\nu=0$
@@
Thus, we have
$\dfrac{1}{2}s_{l}\dfrac{\partial\hat{L}}{\partial s_{l}}+k_{l}\dfrac{\partial\hat{L}}{\partial k_{l}}=k_{l}s_{l}\Bigl{(}s_{l}\log e-\tilde{\mu}_{0}(p)\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl{)}-\nu k_{l}=0.$
@@
For all $1\leq l\leq 3$ it holds $s_{l}^{2}\log e=\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl{)}+\nu$, which can take place, if the set $\{s_{1},s_{2},s_{3}\}$ has no more than two different values (which is realized, formally, for fractional multiplicities). This can be regarded as the next case.
@@
Up to permutation and the global sign, $(z_{j})_{j=1}^{m}=(\underbrace{s_{1},...,s_{1}}_{k_{1}\text{ times}},\underbrace{s_{2},...,s_{2}}_{k_{2}\text{ times}})$, $s_{1}>0$}, $s_{2}<0$}. The components of the minimizer should have different signs; otherwise, changing the sign of a single component would yield a lower value for $\tilde{F}$.